Please help me answer this question:
Which number is 1/7 of the way from 1/4 to 3/5?
F. 1/28
G. 11/28
H. 7/20
J. 3/10
K. 2/7
Thank You very MUCH!!!

Answer by jim_thompson5910(35256)

First find the difference between the two numbers 1/4 and 3/5

3/5 - 1/4 = 12/20 - 5/20 = 7/20

So this distance is 7/20

Now divide this distance by 7

(7/20)/7 = (7/20)*(1/7) = 7/140 = 1/20

So one-seventh of the distance from 1/4 to 3/5 is 1/20.

Finally, add this to the first number (the smaller number) 1/4 to get

1/4 + 1/20 = 5/20 + 1/20 = 6/20 = 3/10

So the answer is choice J